Solve the equation

(a) 5(2 − x) ≤ 3x + 2 < 8

(b) 1/8 = 23x−4,

(c) logbx + logb(x − 4) = logb21

(d) log42x = 2



1
Expert's answer
2022-02-21T16:27:36-0500

a)

"10-5x\le3x+2"

"8\le8x"

"1\le\>x"

"3x+2<8"

"3x<6"

"x<2"

"1\le\>x<2"


b)


"2^{-3}=2^{3x-4}"

"-3=3x-4"

"1=3x"

"x=\frac{1}{3}"



c)


Log"_b" "\begin{bmatrix}
 x(x-4)
 
\end{bmatrix}=" log"_b21"

"x^2-4x-21=0"

"x^2-7x+3x-21=0\\
x(x-7)+3(x-7)=0\\(x+3)(x-7)=0\\
x=-3\>or\>7"


d)


"2x=4^2"

"x=8"